OpenStudy (anonymous):

perform the indicated operation: (a+3)^2/a-3 ÷ 5/(a^2-9)

OpenStudy (zarkon):

\[\frac{(a+3)^3}{a-3}\frac{a^2-9}{5}\] \[=\frac{(a+3)^3}{a-3}\frac{(a-3)(a+3)}{5}\] \[=\frac{(a+3)^4}{5}\]